# Potato Potahto in the FAO-GAEZ Productivity Measures?

### v3

@@ Abstract
The FAO-GAEZ productivity data are widely used in Economics. However, the empirical literature rarely discusses measurement error. We use two proxies to derive novel bounds around the effect of agricultural productivity in a setting with nonclassical measurement error. These bounds rely on assumptions that are weaker than the ones imposed in empirical studies and exhaust the information contained in the first two moments of the data. We reevaluate three influential studies, documenting that measurement error matters and that the impact of agricultural productivity may be smaller than previously reported. Our methodology has broad applications in empirical research involving mismeasured variables.
@@ Introduction
To address this issue, we propose a novel econometric method that uses multiple proxies to partially identify the effect of agricultural productivity, accounting for measurement error. To illustrate the empirical relevance of our approach, we revisit three influential studies: Nunn and Qian (2011), Bustos et al. (2016), and Acharya et al. (2016). Applying our method with two available proxies, we find that the estimated effect of agricultural productivity on population might be substantially smaller than reported by Nunn and Qian (2011); that there might be no effect on political behavior, in contrast to the findings of Acharya et al. (2016); and that the effect on structural transformation remains robust, consistent with Bustos et al. (2016).
@@
Motivated by our assessment of the empirical studies using the FAO-GAEZ productivity measures, we propose to partially identify the effect of agricultural productivity in a linear model with two proxies that are subject to nonclassical measurement error. In this linear model, we impose that the true unobservable productivity is exogenous and that its effect on the outcome is non-negative. [^1] Knowledge of the direction of the coefficient of interest may be motivated by economic theory or by assuming that measurement error issues are not severe enough to change the sign of the estimated coefficient in the empirical literature. If a researcher is not willing to impose this type of assumption, Appendix D proposes an alternative way to partially identify this coefficient without making assumptions about its sign. We also impose that the measurement errors in the two productivity proxies are exogenous in the sense that they are uncorrelated with the economic shock of the outcome variable. Importantly, we allow these measurement errors to be nonclassical and only restrict them to be positively correlated with each other. [^2] Positive correlation between the measurement errors in the two proxies is motivated by arguing that unobservable natural productivity factors in the FAO-GAEZ version 4 model are also unobservable in the FAO-GAEZ version 3 model. If a researcher is not willing to impose this type of assumption, Appendix C proposes an alternative way to partially identify this coefficient without restricting the correlation between the two measurement error terms.
Under these assumptions, we derive bounds around the true linear effect of agricultural productivity on the outcome of interest. Importantly, our assumptions are no stronger than those commonly applied when ignoring measurement error, and our bounds exhaust all the information contained in the first two moments of the data distribution. Intuitively, our bounds create the shortest possible interval that contains the true target parameter and is constructed through restrictions on only the first two moments of the latent variables. Importantly, we focus on this type of restriction because most applied researchers use linear regression models and do not restrict higher-order moments.
Moreover, our bounds have intuitive closed-form solutions. First, the upper bound captures an extreme scenario when the proxy variables are uninformative about the true productivity, i.e., they contain only noise. Second, the lower bound collapses to the standard OLS estimand when there is no measurement error. As a consequence, the similarity between these two objects may be used as an indirect test of whether measurement error is a relevant issue in empirical applications.
@@
Related Literature. This article contributes to two different branches of literature. Concerning its empirical contribution, many papers in economics have used the FAO-GAEZ data. Nunn and Qian (2011), in pioneering work, uses the potential productivity of potatoes to explain how its introduction in the Old World affected population and urbanization. Bustos et al. (2016, 2019, 2020) use the productivity of soybeans to measure the impact of introducing genetically modified seeds on structural transformation. Bentzen et al. (2017) uses productivity under rain-fed and irrigation conditions to study the impacts of irrigation on the origins of autocracies. Acharya et al. (2016) uses cotton productivity as an instrument for slavery to explain contemporary differences in political attitudes in the U.S. Carillo (2021) uses wheat productivity to study the impact of agricultural policies on industrialization in Italy. [^3] Similar papers include Moorthy (2022), Dias et al. (2023) and Carreira et al. (2024) on soybean productivity and gender disparity, water pollution, and deforestation; Galor and Özak (2016) on the productivity of crops available before and after the Columbian Exchange and time preferences; Alesina et al. (2011, 2013) on the productivity of various crops, plough use, and gender disparities; Dube et al. (2016) on maize productivity and the Mexican drug sector; Gollin et al. (2021) on the productivity of various crops and the impacts of the Green Revolution; Mayshar et al. (2022) on the relative productivity of cereals over roots and tubers and the origin of states; Fiszbein (2022) on the productivity of multiple crops and its effects on development. These papers use FAO-GAEZ data to estimate linear specifications of the impacts of crop productivity on various outcomes of interest. However, it is worth noting that there is a range of papers employing FAO-GAEZ data in different types of models, such as trade models (Costinot et al., 2016; Sotelo, 2020; Pellegrina, 2022; Farrokhi and Pellegrina, 2023), deforestation models (Sant’Anna, 2024; Araujo et al., 2025; Dominguez-Iino, 2021), and other types of spatial models (Adamopoulos and Restuccia, 2022; Adamopoulos et al., 2022; Imbert et al., 2022; Fernández-Villaverde et al., 2023).
@@
Concerning its theoretical contribution, our work is inserted in the literature about identifying treatment effect parameters with measurement error. As illustrated by Schennach (2016) and Hu (2017), this literature is vast and growing. The closest paper to ours was written by Lubotsky and Wittenberg (2006), who also focus on a linear model with two mismeasured proxies whose measurement error terms are possibly correlated with each other. However, differently from our approach, they impose that the measurement error terms are uncorrelated with the true variable of interest.
We also highlight that our work is connected to the studies developed by Hausman et al. (1998), Black et al. (2000), Cunha et al. (2010), Alix-García and Millimet (2023) and Millimet and Parmeter (2024). Similarly to us, Black et al. (2000) and Millimet and Parmeter (2024) focus on nonclassical measurement error problems. However, differently from our methodology, they impose that the true variable of interest and the measurement error are negatively correlated with each other. Furthermore, Hausman et al. (1998) and Alix-García and Millimet (2023) focus on mismeasured binary outcome variables while our model analyzes a mismeasured independent variable. Lastly, Cunha et al. (2010) require three proxies to achieve point-identification in a model with nonclassical measurement error, while our approach requires only two proxies to partially identify the coefficient of interest.
@@
There is no precise documentation of the differences between versions. Upon inspection of versions 3 and 4, we identified many differences in parameters and raw data. To illustrate this issue, we provide four examples. First, the “harvest index” parameter essentially measures the portion of the weight of the crop that is actually of economic interest — e.g., the beans of soybeans — and is different for a set of crops between versions. Second, for another set of crops, FAO-GAEZ updated the “leaf area index”, which measures the plant’s canopy and its ability to capture sunlight. Third, precipitation data in version 3 is from the Variability Analysis of Surface Climate Observations 1.1 (VASClimO), while version 4’s precipitation data is from the Global Precipitation Climatology Centre (GPCC) Full Data Reanalysis Product Version 6. Fourth, data that still uses the same source are using newer versions. For example, temperature, sunshine fraction, and wind speed were from CRU 1.1 in version 3 and from CRU 3.21 in version 4, which entails a range of internal modifications of each data source. [^5] Rhode (2024) discusses differences in the model for the productivity of cotton. Despite these differences, all versions of the FAO-GAEZ intend to measure the same potential productivity, i.e., the agricultural productivity for a given place, climate, technology, and water supply.
@@
To illustrate the importance of measurement error in FAO-GAEZ data and motivate our econometric framework in the next section, we highlight three applications in the literature: Nunn and Qian (2011), Bustos et al. (2016), Acharya et al. (2016). In the remainder of this section, we discuss some of the differences in agricultural productivity measured by different versions of FAO-GAEZ for the crops studied in these three articles.
@@
where $Y_{res}\coloneqq Y-W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot Y\right]$, $X_{res}^{*}\coloneqq X^{*}-W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot X^{*}\right]$, $\epsilon_{res}\coloneqq\epsilon-W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot\epsilon\right]$, and $W=\left[1,W_{1},\ldots,W_{J}\right]$ is the row vector of covariates including the constant term.
@@
enforcing Assumption 2.
@@
Assumption 3 imposes that the effect of the true variable of interest on the outcome variable is non-negative. [^7] The case that imposes that the true target parameter is non-positive is symmetric. Moreover, if the researcher is unwilling to impose Assumption 3, Appendix D offers an alternative partial identification strategy that does not require prior knowledge of the direction of the true effect. As a consequence of imposing fewer assumptions, the identified interval around the parameter of interest is weakly wider. Importantly, this assumption does not rule out that the effect might be zero. In our empirical application, it imposes that the effect of the true crop productivity on the outcome variable is non-negative. This assumption is motivated by the empirical literature because we assume that the articles using one proxy — the FAO-GAEZ version 3 — are at least identifying the correct sign of the true effect. If a researcher is not willing to impose this type of assumption, Appendix D proposes an alternative way to partially identify this coefficient without making assumptions about its sign.
@@
In our empirical application, Assumption 4 intuitively imposes that prediction errors in the FAO-GAEZ measures are uncorrelated with the unobserved economic forces that determine the outcome variable. This assumption would be violated if the institutional factors that impact $\epsilon_{res}$ positively also affect the quality of the environmental data that is used when constructing the FAO-GAEZ measure. In this case, regions with large values of $\epsilon_{res}$ would have a small amount of measurement error. Consequently, it is implausible that $Cov\left(U_{1},\epsilon_{res}\right)=0$.
@@
Assumption 5 imposes that the covariance between the measurement error terms is positive. In our empirical application, it intuitively imposes that areas with large prediction errors in the FAO-GAEZ version 3 model tend to have large prediction errors in the FAO-GAEZ version 4 model too. This assumption is plausible because unobservable natural productivity factors in the FAO-GAEZ version 4 model are also unobservable in the FAO-GAEZ version 3 model. [^9] Appendix C offers an alternative partial identification strategy that does not restrict the covariance between the two measurement error terms. As a consequence of imposing fewer assumptions, the identified interval around the parameter of interest is weakly wider.
@@
Moreover, the bounds in Equation (6) exhaust all the information contained in the first two moments of the observable variables $\left(Y_{res},Z_{1,res},Z_{2,res}\right)$.
@@
First, the lower bound in Equation (4) collapses to the standard OLS estimand when there is no measurement error. Consequently, similarity between the estimated lower bound and the estimated OLS coefficient may be used as an indirect test of whether measurement error is a relevant issue in a given empirical application.
@@
$y_{jt}=\alpha_{j}+\alpha_{t}+\beta\cdot I^{Post}\cdot\log(1+A_{j}^{potato})+\sum_{l=1}^{L}\alpha_{l}\cdot W_{l,j}+\epsilon_{jt},$ (8)
where $y_{jt}$ is population in country $j$ and year $t$, $A_{j}^{potato}$ is the area of country $j$ that is suitable for the cultivation of white potato, [^12] A pixel is defined as being suitable if it has $\geq 40\%$ of the maximum attainable yield for the crop. $I^{Post}$ is a dummy variable that equals one for years after 1700, i.e., after the introduction of the potato. In our setting, the interaction $I^{Post}\cdot\log(1+A_{j}^{potato})$ is the variable of interest $X^{*}$.
@@
Table 1 shows descriptive statistics for the area suitable for potatoes as measured by the FAO-GAEZ proxies. In version 4, the mean, standard deviation and interquantile range of suitable area increases. Consequently, there is more spatial heterogeneity in version 4 than in version 3.
Table 1: Descriptive Statistics: Area Suitable for Potato
$1+A_{j}^{potato}$ $\log(1+A_{j}^{potato})$ $\Delta\log(1+A_{j}^{potato})$
V 3 V 4 V 3 V 4 V 4 - V 3
(1) (2) (3) (4) (5)
Mean 4,646 8,583 3.88 4.63 0.75
S.D. 22,607 37,355 3.91 4.17 1.40
Min 1 1 0 0 -1.95
1st Quartile 1 1 0 0 0
Median 61 353 4.10 5.86 0.07
3rd Quartile 2,022 5,283 7.61 8.57 0.94
Max 248,816 403,667 12.42 12.90 7.10
Notes: This table shows descriptive statistics of potato suitable areas for FAO-GAEZ versions 3 and 4 with the low input specification. Columns (1) and (2) are measured in thousands of hectares. The unit of observation is the country-year pair. “S.D.” stands for standard deviation.
@@
Notes: Observations are at the country-year level. All regressions use a baseline sample of 130 Old World countries. Countries in North and South America are excluded. The periods are 1000, 1100, 1200, 1300, 1400, 1500, 1600, 1700, 1750, 1800, 1850, and 1900. Column (1) reports the result of Nunn and Qian (2011)’s table IV column 5 which uses middle input potato suitability from FAO-GAEZ version 2. To make results more comparable, Columns (2) and (3) report the results for using low input FAO GAEZ data version 3 and 4, respectively. Columns (4) and (5) show the result when $\log(1+A_{j}^{potato})$ is normalized to have a variance equal to one. Column (6) presents the result of our partial identification approach. All specifications include year-fixed effects, country-fixed effects, and a range of controls in logs interacted with year dummies: Old World Crops Area, Elevation, Ruggedness, Tropical Area, Maize Area, Silage Maize Area, Sweet Potatoes Area, Cassava Area. Standard errors are clustered at the country level. The confidence interval in Column (6) is built at the 10% and 1% levels.
@@
We can now discuss the results of our partial identification estimator. The interpretation of these results does not align directly with the results shown so far for two reasons. First, in our approach, we normalize the variance of $\log(1+A_{j}^{potato})$ such that it equals one (Assumption 2). Second, Nunn and Qian (2011) uses the transformation $\log(1+x)$ which makes it hard to interpret the estimated coefficient as elasticities (Chen and Roth, 2023). For these two reasons, we discuss our results in terms of standard deviations of $\log(1+A_{j}^{potato})$. To make the comparison easier, Columns (4) and (5) show the result of estimating Equation (8) with the normalized measures from FAO-GAEZ versions 3 and 4. In Columns (4) and (5), the interpretation of the results is that an increase of one standard deviation in $\log(1+A_{j}^{potato})$ causes an increase of 4.3% and 3.8% in population.
Column (6) shows the results associated with our method. As we develop a partial identification strategy, our result is an estimated interval. This interval shows a wide range of possible effects, with an increase of one standard deviation in $\log(1+A_{j}^{potato})$ leading to an increase in population between 2.9% and 22.9%. Importantly, this interval exhausts all the information contained in the first and second moments of the data-generating process, implying that the large estimate for the upper bound is not a result of our estimator being loose. Such a wide estimated interval, especially in comparison to the small standard errors reported in Columns (4) and (5), suggests that model uncertainty with respect to the existence of non-classical measurement error is more concerning than sample uncertainty.
We also report in Column (6) the confidence intervals at the 10% and 1% levels. Even at the 10%-confidence level, we see that the lower bound of the effect is a quarter of the point-estimate associated with the FAO GAEZ version 3 (Column (4)). We also highlight that, when ignoring measurmente error, Nunn and Qian (2011)’s results and our results in Columns (2)-(5) are all significantly different from zero at the 1%-confidence level. However, when we account for measurement error (Column (6)), we cannot reject that the effect of the introduction of potato in population is zero at the 1%-confidence level.

### v4

@@ Abstract
The FAO-GAEZ productivity data are widely used in Economics. However, the empirical literature rarely discusses measurement error. We use two proxies to derive novel analytical bounds around the effect of agricultural productivity in a setting with nonclassical measurement error. These bounds rely on assumptions that are weaker than the ones imposed in empirical studies and exhaust the information contained in the first two moments of the data. We reevaluate three influential studies, documenting that measurement error matters and that the impact of agricultural productivity may be smaller than previously reported. Our methodology has broad applications in empirical research involving mismeasured variables.
@@ Introduction
To address this issue, we propose a novel analytical method that uses multiple proxies to partially identify the effect of agricultural productivity, accounting for measurement error. To illustrate the empirical relevance of our approach, we revisit three influential studies: Nunn and Qian (2011), Bustos et al. (2016), and Acharya et al. (2016). Applying our method with two available proxies, we find that the estimated effect of agricultural productivity on population might be substantially smaller than reported by Nunn and Qian (2011); that there might be no effect on political behavior, in contrast to the findings of Acharya et al. (2016); and that the effect on structural transformation remains robust, consistent with Bustos et al. (2016).
@@
Motivated by our assessment of the empirical studies using the FAO-GAEZ productivity measures, we propose to partially identify the effect of agricultural productivity in a linear model with two proxies that are subject to nonclassical measurement error. In this linear model, we impose that the true unobservable productivity is exogenous and that its effect on the outcome is non-negative. [^1] Knowledge of the direction of the coefficient of interest may be motivated by economic theory or by assuming that measurement error issues are not severe enough to change the sign of the estimated coefficient in the empirical literature. If a researcher is not willing to impose this type of assumption, Appendix D proposes an alternative way to partially identify this coefficient without making assumptions about its sign. We also impose that the measurement errors in the two productivity proxies are exogenous in the sense that they are uncorrelated with the economic shock of the outcome variable. Importantly, we allow these measurement errors to be nonclassical and only restrict them to be positively correlated with each other. [^2] Positive correlation between the measurement errors in the two proxies is motivated by arguing that unobservable natural productivity factors in the FAO-GAEZ version 4 model are also unobservable in the FAO-GAEZ version 3 model (Appendix I). If a researcher is not willing to impose this type of assumption, Appendix C proposes an alternative way to partially identify this coefficient without restricting the correlation between the two measurement error terms.
Under these assumptions, we derive analytical bounds around the true linear effect of agricultural productivity on the outcome of interest. Importantly, our assumptions are no stronger than those commonly applied when ignoring measurement error, and our bounds exhaust all the information contained in the first two moments of the data distribution. Intuitively, our bounds create the shortest possible interval that contains the true target parameter and is constructed through restrictions on only the first two moments of the latent variables. Importantly, we focus on this type of restriction because most applied researchers use linear regression models and do not restrict higher-order moments.
Moreover, our bounds have intuitive closed-form solutions. First, the upper bound captures an extreme scenario when the proxy variables are uninformative about the true productivity, i.e., they contain only noise. Second, the lower bound collapses to the standard OLS estimand when there is no measurement error. As a consequence of this necessary condition, the similarity between these two objects may be used as a suggestive assessment of whether measurement error is a relevant issue in empirical applications.
@@
Related Literature. This article contributes to two different branches of literature. Concerning its empirical contribution, many papers in economics have used the FAO-GAEZ data. Nunn and Qian (2011), in pioneering work, uses the potential productivity of potatoes to explain how its introduction in the Old World affected population and urbanization. Bustos et al. (2016); Bustos et al. (2019); Bustos et al. (2020) use the productivity of soybeans to measure the impact of introducing genetically modified seeds on structural transformation. Bentzen et al. (2017) uses productivity under rain-fed and irrigation conditions to study the impacts of irrigation on the origins of autocracies. Acharya et al. (2016) uses cotton productivity as an instrument for slavery to explain contemporary differences in political attitudes in the U.S. Carillo (2021) uses wheat productivity to study the impact of agricultural policies on industrialization in Italy. [^3] Similar papers include Moorthy (2022), Dias et al. (2023) and Carreira et al. (2024) on soybean productivity and gender disparity, water pollution, and deforestation; Galor and Özak (2016) on the productivity of crops available before and after the Columbian Exchange and time preferences; Alesina et al. (2011, 2013) on the productivity of various crops, plough use, and gender disparities; Dube et al. (2016) on maize productivity and the Mexican drug sector; Gollin et al. (2021) on the productivity of various crops and the impacts of the Green Revolution; Mayshar et al. (2022) on the relative productivity of cereals over roots and tubers and the origin of states; Fiszbein (2022) on the productivity of multiple crops and its effects on development. These papers use FAO-GAEZ data to estimate linear specifications of the impacts of crop productivity on various outcomes of interest. However, it is worth noting that there is a range of papers employing FAO-GAEZ data in different types of models, such as trade models (Costinot et al., 2016; Sotelo, 2020; Pellegrina, 2022; Farrokhi and Pellegrina, 2023), deforestation models (Sant’Anna, 2024; Araujo et al., 2025; Dominguez-Iino, 2021), and other types of spatial models (Adamopoulos and Restuccia, 2022; Adamopoulos et al., 2022; Imbert et al., 2022; Fernández-Villaverde et al., 2023).
@@
Concerning its theoretical contribution, our work is inserted in the literature about identifying treatment effect parameters with measurement error. As illustrated by Schennach (2016) and Hu (2017), this literature is vast and growing. The closest paper to ours was written by Lubotsky and Wittenberg (2006). While we derive an identified set that contains the true target parameter, they propose a single estimand that minimizes bias. Moreover, they impose that the measurement error terms are uncorrelated with the true variable of interest, while we impose no restriction on this correlation. This extra flexibility is particularly important in our empirical application because measurement error terms must be correlated with the true variable of interest when the latter is bounded, as is the case with crop productivity.
We also highlight that our work is connected to the studies developed by Hausman et al. (1998), Black et al. (2000), Hu and Schennach (2008), Cunha et al. (2010), Schennach (2014), Hu and Sasaki (2015), Alix-García and Millimet (2023) and Millimet and Parmeter (2024). Black et al. (2000) and Millimet and Parmeter (2024) differ from our work because they impose that the true variable of interest and the measurement error must be negatively correlated with each other. Furthermore, Hausman et al. (1998) and Alix-García and Millimet (2023) focus on mismeasured binary outcome variables. Moreover, Hu and Schennach (2008), Cunha et al. (2010) and Hu and Sasaki (2015) require the measurement error terms to be mutually independent. Lastly, Schennach (2014) proposes a general method that encompasses our setting. However, her tool requires simulation-based optimization algorithms while our tool is based on an analytic solution that is easily implemented using standard statistical software. The importance of analytical solutions such as ours is emphasized by Schennach (2014, page 366).
@@
There is no precise documentation of the differences between versions. Upon inspection of versions 3 and 4, we identified many differences in parameters and raw data. To illustrate this issue, we provide four examples. First, the “harvest index” parameter essentially measures the portion of the weight of the crop that is actually of economic interest — e.g., the beans of soybeans — and is different for a set of crops between versions. Second, for another set of crops, FAO-GAEZ updated the “leaf area index”, which measures the plant’s canopy and its ability to capture sunlight. Third, precipitation data in version 3 is from the Variability Analysis of Surface Climate Observations 1.1 (VASClimO), while version 4’s precipitation data is from the Global Precipitation Climatology Centre (GPCC) Full Data Reanalysis Product Version 6. Fourth, data that still uses the same source are using newer versions. For example, temperature, sunshine fraction, and wind speed were from CRU 1.1 in version 3 and from CRU 3.21 in version 4, which entails a range of internal modifications of each data source. [^5] Rhode (2024) discusses differences in the model for the productivity of cotton. Despite these differences, all versions of the FAO-GAEZ intend to measure the same potential productivity, i.e., the agricultural productivity for a given place, climate, technology, and water supply. Appendix I further discusses sources of measurement error in FAO-GAEZ data.
@@
To illustrate the importance of measurement error in FAO-GAEZ data and motivate our econometric framework in the next section, we highlight three applications in the literature: Nunn and Qian (2011), Bustos et al. (2016), Acharya et al. (2016). In the remainder of this section, we discuss some of the differences in agricultural productivity measured by different versions of FAO-GAEZ for the crops studied in these three articles. Importantly, the productivity measures used by Nunn and Qian (2011) and Acharya et al. (2016) are bounded, implying that any measurement error they may contain must be nonclassical.
@@
where $Y_{res}\coloneqq Y-W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot Y\right]$, $X_{res}^{*}\coloneqq X^{*}-W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot X^{*}\right]$, $\epsilon_{res}\coloneqq\epsilon-W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot\epsilon\right]=\epsilon$, and $W=\left[1,W_{1},\ldots,W_{J}\right]$ is the row vector of covariates including the constant term.
@@
After partialling out the covariate vector $W$, the true unobserved measure $X^{*}$ is exogenous, i.e., $Cov\left(X_{res}^{*},\epsilon_{res}\right)=0$.
@@
enforcing Assumption 2. One may view this normalization as changing the target parameter from $\tilde{\beta}$ to $\beta=\tilde{\beta}\cdot\sqrt{Var\left(\tilde{X}_{res}^{*}\right)}$ since it is not possible to separately identify $\tilde{\beta}$ and $\sqrt{Var\left(\tilde{X}_{res}^{*}\right)}$.
@@
Assumption 3 imposes that the effect of the true variable of interest on the outcome variable is non-negative. [^7] The case that imposes that the true target parameter is non-positive is symmetric. Importantly, this assumption does not rule out that the effect might be zero. In our empirical application, it imposes that the effect of the true crop productivity on the outcome variable is non-negative. This assumption is motivated by the empirical literature because we assume that the articles using one proxy — the FAO-GAEZ version 3 — are at least identifying the correct sign of the true effect. If a researcher is not willing to impose this type of assumption, Appendix D proposes an alternative way to partially identify this coefficient without making assumptions about its sign.
@@
In our empirical application, Assumption 4 intuitively imposes that prediction errors in the FAO-GAEZ measures are uncorrelated with the unobserved economic forces that determine the outcome variable. This assumption would be violated if the institutional factors that impact $\epsilon_{res}$ positively also affect the quality of the environmental data that is used when constructing the FAO-GAEZ measure. In this case, regions with large values of $\epsilon_{res}$ would have a small amount of measurement error. Consequently, $Cov\left(U_{1},\epsilon_{res}\right)=0$ would be implausible in this hypothetical scenario.
@@
Assumption 5 imposes that the covariance between the measurement error terms is positive. In our empirical application, it intuitively imposes that areas with large prediction errors in the FAO-GAEZ version 3 model tend to have large prediction errors in the FAO-GAEZ version 4 model too. This assumption is plausible because unobservable natural productivity factors in the FAO-GAEZ version 4 model are also unobservable in the FAO-GAEZ version 3 model. [^9] Appendix C offers an alternative partial identification strategy that does not restrict the covariance between the two measurement error terms. As a consequence of imposing fewer assumptions, the identified interval around the parameter of interest is weakly wider. In Appendix I, we discuss sufficient conditions in the FAO-GAEZ data under which this assumption holds.
@@
Moreover, the bounds in Equation (6) exhaust all the information contained in the first two moments of the observable variables $\left(Y_{res},Z_{1,res},Z_{2,res}\right)$. [^11] See Appendix A.1.4 for the formal definition of “exhausting all the information contained in the first two moments of the observable variables.” Below, we provide a intuitive definition for this concept.
@@
First, the lower bound in Equation (4) collapses to the standard OLS estimand when there is no measurement error. Consequently, similarity between the estimated lower bound and the estimated OLS coefficient may be used as an indirect test of whether measurement error is a relevant issue in a given empirical application. Importantly, this test relies on a necessary condition of “no measurement error” instead of a sufficient condition. For this reason, it should be used as a suggestive assessment only, and it should be combined with an economic argument that explains why measurement error is irrelevant in a specific application.
@@
Remark 1
If a researcher is willing to impose more assumptions on the data-generating process, she may tighten the bounds derived in Proposition 1. For example, imagine that the second proxy, $Z_{2,res}$, is better than the first proxy, $Z_{1,res}$, in the sense that $Var\left(U_{1}\right)\geq Var\left(U_{2}\right)$. Under this additional assumption, we tighten the bounds in Equation (6) with the following restriction:
$\left\{\begin{array}[]{rclcl}\beta&\geq&2\cdot\left[\frac{Cov\left(Z_{1,res},Y_{res}\right)-Cov\left(Z_{2,res},Y_{res}\right)}{Var\left(Z_{1,res}\right)-Var\left(Z_{2,res}\right)}\right]&\text{if}&Var\left(Z_{1,res}\right)>Var\left(Z_{2,res}\right)\\ \beta&\leq&2\cdot\left[\frac{Cov\left(Z_{1,res},Y_{res}\right)-Cov\left(Z_{2,res},Y_{res}\right)}{Var\left(Z_{1,res}\right)-Var\left(Z_{2,res}\right)}\right]&\text{if}&Var\left(Z_{1,res}\right)<Var\left(Z_{2,res}\right)\end{array}\right.$.
@@
To map our assumptions to the three papers, we note the following. Assumption 1 (Exogeneity) is imposed in all of them. Assumption 2 (Scale Normalization) only affects the interpretation of results, which are expressed in terms of standard deviations of $X^{*}$. Assumption 3 (Known Direction of the True Effect) requires that each application correctly identifies at least the sign of the effect. [^13] Appendix D shows how this assumption can be relaxed in our setting. Assumption 4 (Exogeneity of Measurement Errors) weakens the identification requirements relative to those in the papers. Assumption 5 (Nonclassical Measurement Error) is motivated by the persistence of measurement error in FAO-GAEZ data and is further discussed in Appendix I. Finally, Assumption 6 (Well-behaved Data Distributions) is a standard technical condition.
@@
$y_{it}=\alpha_{i}+\alpha_{t}+\beta\cdot I^{Post}\cdot\log(1+A_{i}^{potato})+\sum_{j=1}^{J}\alpha_{j}\cdot w_{i,j}+\epsilon_{it},$ (8)
where $y_{it}$ is population in country $i$ and year $t$, $A_{i}^{potato}$ is the area of country $i$ that is suitable for the cultivation of white potato, [^14] A pixel is defined as being suitable if it has $\geq 40\%$ of the maximum attainable yield for the crop. $I^{Post}$ is a dummy variable that equals one for years after 1700, i.e., after the introduction of the potato. In our setting, the interaction $I^{Post}\cdot\log(1+A_{i}^{potato})$ is the variable of interest $X^{*}$.
@@
Table 1 shows descriptive statistics for the interaction between the Post-1700 indicator variable and the natural logarithm of the area suitable for potatoes as measured by the FAO-GAEZ proxies. In version 4, the mean, standard deviation and maximum of this interaction increase. Consequently, there is more spatial heterogeneity in version 4 than in version 3.
Table 1: Descriptive Statistics: Area Suitable for Potato (Interaction)
$I^{Post}\cdot\log(1+A_{i}^{potato})$ $\Delta\left\{I^{Post}\cdot\log(1+A_{i}^{potato})\right\}$
V 3 V 4 V 4 - V 3
(1) (2) (3)
Mean 1.31 1.56 0.25
S.D. 2.92 3.27 0.89
Min 0 0 -1.96
1st Quartile 0 0 0
Median 0 0 0
3rd Quartile 0 0 0
Max 12.42 12.91 7.11
Notes: This table shows descriptive statistics for the interaction between the Post-1700 indicator variable and the natural logarithm of potato suitable areas for FAO-GAEZ versions 3 and 4 with the low input specification. The unit of observation is the country-year pair. “S.D.” stands for standard deviation.
@@
Notes: Observations are at the country-year level. All regressions use a baseline sample of 130 Old World countries. Countries in North and South America are excluded. The periods are 1000, 1100, 1200, 1300, 1400, 1500, 1600, 1700, 1750, 1800, 1850, and 1900. Column (1) reports the result of Nunn and Qian (2011)’s table IV column 5 which uses middle input potato suitability from FAO-GAEZ version 2. To make results more comparable, Columns (2) and (3) report the results for using low input FAO-GAEZ data version 3 and 4, respectively. Columns (4) and (5) show the result when $I^{Post}\cdot\log(1+A_{i}^{potato})$ is normalized to have a variance equal to one. Column (6) presents the result of our partial identification approach. All specifications include year-fixed effects, country-fixed effects, and a range of controls in logs interacted with year dummies: Old World Crops Area, Elevation, Ruggedness, Tropical Area, Maize Area, Silage Maize Area, Sweet Potatoes Area, Cassava Area. Standard errors are clustered at the country level. The confidence interval in Column (6) is built at the 10% and 1% levels.
@@
We can now discuss the results of our partial identification estimator. The interpretation of these results does not align directly with the results shown so far for two reasons. First, in our approach, we normalize the variance of $I^{Post}\cdot\log(1+A_{i}^{potato})$ such that it equals one (Assumption 2). Second, Nunn and Qian (2011) uses the transformation $\log(1+x)$ which makes it hard to interpret the estimated coefficient as elasticities (Chen and Roth, 2023). For these two reasons, we discuss our results in terms of standard deviations of $I^{Post}\cdot\log(1+A_{i}^{potato})$. To make the comparison easier, Columns (4) and (5) show the result of estimating Equation (8) after normalizing $I^{Post}\cdot\log(1+A_{i}^{potato})$ for each FAO-GAEZ version. In Columns (4) and (5), the interpretation of the results is that an increase of one standard deviation in $I^{Post}\cdot\log(1+A_{i}^{potato})$ causes an increase of 4.3% and 3.8% in population.
Column (6) shows the results associated with our method. As we develop a partial identification strategy, our result is an estimated interval. This interval shows a wide range of possible effects, with an increase of one standard deviation in $I^{Post}\cdot\log(1+A_{i}^{potato})$ leading to an increase in population between 2.9% and 22.9%. Importantly, this interval exhausts all the information contained in the first and second moments of the data-generating process, implying that the large estimate for the upper bound is not a result of our estimator being loose. Such a wide estimated interval, especially in comparison to the small standard errors reported in Columns (4) and (5), suggests that model uncertainty with respect to the existence of non-classical measurement error is more concerning than sample uncertainty.
We also report in Column (6) the confidence intervals at the 10% and 1% levels. Even at the 10%-confidence level, we see that the lower bound of the effect is a quarter of the point-estimate associated with the FAO-GAEZ version 3 (Column (4)). We also highlight that, when ignoring measurement error, Nunn and Qian (2011)’s results and our results in Columns (2)-(5) are all significantly different from zero at the 1%-confidence level. However, when we account for measurement error (Column (6)), we cannot reject that the effect of potatoes introduction on population is zero at the 1%-confidence level.